OpenStudy (happykiddo):

The figures below show two different ways of arranging four identical triangles of grey poster board on top of a white square. The square has sides equal to x + y, while the hypotenuse of each triangle is represented by the variable c.

OpenStudy (happykiddo):

Hazel wrote the following statements to prove that c2 = x2 + y2. 1. Area of the four grey triangles inside figure A = 2. Area of the white square inside figure A = c2 3. Area of figure A = c2 + 2xy 4. Area of the four grey triangles inside figure B = 5. Area of the two white squares inside figure B = 2x2+ 2y2 6. Area of figure B = 2x2+ 2y2 + 2xy 7. Area of figure A = area of Figure B, hence c2 + 2xy = 2x2+ 2y2 + 2xy 8. Therefore, c2 = x2+ y2 Which is the first incorrect statement in Hazel’s proof?

OpenStudy (happykiddo):

Statement 4 Statement 5 Statement 6 Statement 7

OpenStudy (happykiddo):

hi !!

OpenStudy (paxpolaris):

Statements 1 and 4 should both say the area of the 4 grey triangles = \(\large 2xy\). 5. is the first incorrect statement. it should be: the area of the two white squares in figure B = \(\Large x^2+y^2\)

OpenStudy (happykiddo):

i thought it would be 5 also thanks :D